# Dance Any Beat: Blending Beats with Visuals in Dance Video Generation

### v2

@@ Abstract
Automated choreography advances by generating dance from music. Current methods create skeleton keypoint sequences, not full dance videos, and cannot make specific individuals dance, limiting their real-world use. These methods also need precise keypoint annotations, making data collection difficult and restricting the use of self-made video datasets. To overcome these challenges, we introduce a novel task: generating dance videos directly from images of individuals guided by music. This task enables the dance generation of specific individuals without requiring keypoint annotations, making it more versatile and applicable to various situations. Our solution, the Dance Any Beat Diffusion model (DabFusion), utilizes a reference image and a music piece to generate dance videos featuring various dance types and choreographies. The music is analyzed by our specially designed music encoder, which identifies essential features including dance style, movement, and rhythm. DabFusion excels in generating dance videos not only for individuals in the training dataset but also for any previously unseen person. This versatility stems from its approach of generating latent optical flow, which contains all necessary motion information to animate any person in the image. We evaluate DabFusion’s performance using the AIST++ dataset, focusing on video quality, audio-video synchronization, and motion-music alignment. We propose a 2D Motion-Music Alignment Score (2D-MM Align), which builds on the Beat Alignment Score to more effectively evaluate motion-music alignment for this new task. Experiments show that our DabFusion establishes a solid baseline for this innovative task. Video results can be found on our project page: https://DabFusion.github.io.
Figure 1: We present a diffusion-based framework DabFusion to dance any beat given any image. With various music styles of AIST++ dataset, our model predicts matching future video frames (128 $\times$ 128) starting from the same reference image.
@@ 1 Introduction
Generating dance from music has great potential for automating choreography and enhancing various applications. Current methods [25, 26, 46, 34] primarily produce keypoint sequences which, while effective, often yield outputs that are not as user-friendly and cannot animate specific individuals, limiting their use in personalized scenarios. These methods also rely heavily on precisely annotated keypoints, complicating data collection and limiting their application to non-standard or self-made video datasets. To address these limitations, we introduce a new and innovative task: generating dance videos directly from images of individuals, guided by accompanying music. This allows for personalized dance videos featuring specific people without needing keypoint annotations, making it more versatile for different scenarios. Our solution, named the Dance Any Beat Diffusion model (DabFusion), takes a reference image paired with a selected piece of music as input to create dance videos featuring various dance types and choreographies. For extracting information from music, we develop a novel music encoder that captures essential features including dance style, movement, and rhythm, which are critical for this task. To accurately extract dance styles, we fine-tune the CLAP [58] model, a text-audio pretraining foundation model, on the AIST++ [26] dataset, which serves as the training dataset for our task. For movement information, we fine-tune the Wav2CLIP [57] model, a visual-audio pretraining foundation model, on the AIST++ dataset. The synchronization of dance movements with the music’s rhythm is essential for realistic dance generation. To achieve this, we employ Librosa [29], a tool for audio signal analysis, to extract beat information from the music. By integrating these elements, our music encoder can generate a comprehensive representation of the music.
@@
DabFusion effectively generates dance videos not just for individuals included in the AIST++ dataset, but also for people it has never encountered before. As shown in Fig. 1, we animate previously unseen individuals by merging their images with a background from the AIST++ dataset, bringing them to life through dance. DabFusion can choreograph dance sequences for previously unseen individuals to various music pieces, producing distinctly different videos. This diversity in output highlights the robust information extraction capability of our music encoder. When generating dance videos using the AIST++ dataset, as illustrated in Fig. 2, DabFusion excels in producing varied dance styles with different dancers from multiple perspectives and diverse initial poses. The capability of DabFusion to generate various dance forms enables it to serve a broad audience, ranging from professional choreographers seeking inspiration to casual users wanting to see themselves or others dance to their favorite tunes. Additionally, this technology can be utilized in entertainment applications, such as music videos and virtual concerts, where personalized and accurately synchronized dance performances can greatly enhance the viewer experience. The versatility of DabFusion arises from its method of generating latent optical flow, which captures all essential motion information needed to animate any person in the image. This technique is widely adopted in the field of image-to-video (I2V) generation [53, 38, 10, 5, 64, 54, 37, 44, 43, 67]. Inspired by recent I2V works [31, 43, 56], our model’s training process is structured into two phases. First, a latent flow auto-encoder is trained unsupervisedly to estimate the latent optical flow between reference and driving frames in a video, aiding in warping the reference frame to generate movement. Then, a U-Net based diffusion model generates latent flows guided by music and a starting image. We conduct a thorough evaluation of DabFusion’s performance using the AIST++ dataset, with particular focus on video quality, audio-video synchronization, and the alignment between motion and music. To more effectively assess motion-music alignment—a crucial aspect of this new task—we introduce a 2D motion-music alignment (2D-MM Align) score, inspired by the Beat Alignment Score [26] for 3D motion-music correlation. This metric evaluates synchronization between motion and music in 2D scenarios, demonstrating the effectiveness of our approach in producing rhythmically aligned dance videos. Our experimental results demonstrate that DabFusion establishes a robust baseline for this innovative task, indicating its potential as a pioneering tool in the field of dance video synthesis.
@@ 2 Related Works
Generating dance sequences from music uniquely intersects motion synthesis [1, 3, 11, 13] and music interpretation [57, 29, 58], aiming to create choreographed movements synchronized with input music. This extends beyond traditional motion synthesis, as choreographed movements are complex to animate. Early research [42, 68, 8] focused on producing 2D dance sequences due to the availability of online dance videos and advances in 2D human pose estimation [4]. However, 2D predictions lack expressiveness and applicability, prompting a shift to 3D dance generation. Recent methods leverage LSTMs [48, 19, 60], GANs [24, 47, 9, 20], and Transformers [16, 26, 46, 34, 63] for 3D motion generation. The AIST++ dataset [26], a rich compilation of 3D motion data, has significantly advanced the field. Notable models include FACT [26], a full-attention-based cross-modal transformer using sequence-to-sequence learning for lifelike 3D dance sequences, and Bailando [46], which combines a pose VQ-VAE with a Motion GPT for temporal coherence via actor-critic learning. While existing research focuses on generating motion sequences from music, Ren et al. [36] synthesizes videos using these sequences. Our work bypasses intermediary joint sequence generation, directly creating dance videos from music.
Image-to-Video (I2V) generation [2, 7, 15, 28, 33, 53, 59, 61, 65, 54] transforms static images into dynamic video sequences. This domain relates closely to single-image video prediction [21, 27], which infuses still images with temporal continuity to create believable motion. A specialized subfield, conditional image-to-video generation [2, 7, 15, 28, 53, 61, 54], incorporates guiding conditions or cues into the video generation process, enriching the content with specific characteristics or actions. For example, Wang et al. [53] developed a conditional GAN architecture to create human videos from static images, conditioned on labels describing facial expressions or actions. Yang et al. [61] introduced a pose-guided approach that extracts poses from an image and uses GANs to generate a sequence of poses and then video frames, translating static postures into animated sequences. Building on these, Wang et al. [54] unveiled the Latent Image Animator (LIA), which animates still images into videos under the guidance of an auxiliary video, leveraging a self-supervised auto-encoder to navigate latent space for animation. However, using audio, particularly music, as a conditioning factor in I2V remains largely unexplored. Introducing musical elements as conditions for video generation could open new research avenues and applications, suggesting an exciting direction for future work in I2V technology.
Figure 3: Overview of DabFusion. DabFusion incorporates noise input along with image embedding $z_{0}$ and music embedding $e$ as conditions. Following the denoising stage of the diffusion model, we obtain $h_{1}^{N}$, comprising a concatenated sequence of latent flow and corresponding occlusion maps. $h_{1}^{N}$ is utilized to transform $z_{0}$ into a new sequence of latent maps, denoted as $\tilde{z}_{1}^{N}$, which is subsequently decoded to produce an image sequence.
@@ 3.1 Music Encoding
Extracting meaningful information from music is essential for generating dance videos, where dance type, dancer movement, and rhythm play critical roles. For dance type information, we utilize the CLAP model, a pipeline for contrastive language-audio pretraining that develops audio representations by integrating audio data with natural language descriptions. We fine-tune CLAP to enhance its capture of dance style information. Specifically, we freeze the weights of the audio and text encoders in CLAP and introduce multi-layer perceptrons as adapter layers. These layers are trained using data from the AIST++ dataset. For movement information, we employ Wav2CLIP, initially trained on audio-visual datasets. We apply a similar fine-tuning approach as with CLAP, using the AIST++ dataset to adapt Wav2CLIP for our specific requirements. Regarding rhythm, which is crucial for aligning dance movements with the music beat, we use Librosa, a tool for audio signal analysis, to extract beat information. We then concatenate these diverse sources of information to create a comprehensive musical representation tailored for dance video generation.
@@ 4.1 Dataset and Implementation Details
We organize our dataset by music pieces, allocating 10,564 videos and 50 music tracks to the training set, and 2,106 videos with 10 tracks to the test set. Each video is segmented into frames, resized to $128\times 128$ pixels. In the latent flow estimation phase, we adopt the architecture from [18] for both our image encoder and decoder. The flow predictor follows [43]. The model is trained over 150 epochs with a batch size of 100, using the Adam optimizer [22], starting with a learning rate of $2\times 10^{-4}$, decreased by a factor of 0.1 after epochs 60, 90, and 120. For the denoising model’s training, we employ a conditional 3D U-Net architecture from [12], featuring four down-sampling and up-sampling 3D convolutional blocks. The time step $t$ is encoded by the sinusoidal position embedding [50]. The training spanned 250 epochs, starting with the same initial learning rate as the first stage, and adjusted at epochs 100, 150, and 200. We set $T$ = 1000 and apply a cosine noise schedule [32] and dynamic thresholding [40] at the 90% during sampling.
@@ 4.2 Evaluation Metrics
Table 1: Quantitative results between DabFusion and MM-Diffusion.
[…]
MM-Diffusion 180.31 0.025 34.81 0.951
DabFusion (ours) 193.98 0.033 26.88 0.959
@@ 4.3 Result Analysis
Video Quality Evaluation. To evaluate the quality of the videos generated by our DabFusion, we use the music and the first frame from the test set, resulting in 2,106 generated videos. We visually present outputs from our model alongside ground truth in Fig. 5. As shown in Fig. 5, the quality of the videos generated by our model closely approximates that of real videos. Given the novelty of our task and the lack of directly comparable previous works, we benchmark our DabFusion against MM-Diffusion [39], the leading model in unconditional video generation. We select MM-Diffusion because it is trained on a subset of the AIST++ dataset, closely aligning with our training dataset and demonstrating state-of-the-art video quality performance on AIST++. We randomly generate an equivalent number of videos with MM-Diffusion for comparison. A quantitative comparison between our models and MM-Diffusion is provided in Tab. 1. The results reveal that the quality of videos produced by our models closely matches that of MM-Diffusion. These comparisons underscore the competitive quality of videos generated by our approach.
@@
A notable characteristic of the AIST++ dance dataset is its use of nine cameras C01-C09 to capture dance performances, offering multiple perspectives. We analyze four specific angles, summarized in Tab. 2, to assess their impact on video quality. MM-Diffusion, trained on a subset of the AIST++ dataset using only data from the C01 camera, is used to generate an equivalent number of videos for direct comparison. Our models outperform MM-Diffusion in terms of FVD and SSIM scores, especially for camera position C01, demonstrating superior video quality. Analysis of scores among different cameras indicates that direct camera angles, such as C01 and C03, where the cameras point straight at the dancer, result in higher-quality video generation compared to oblique angles like C02. Additionally, we explore the impact of camera distance by comparing videos from C01 and C09, which share the same angle but differ in their proximity to the dancer. The findings suggest that cameras positioned further from the dancer tend to yield better results.
@@ 4.4 Choreograph Anyone
Table 7: Ablation study for result comparison about arbitrary length video generation. We use the last frame of the first generated videos as the starting frame of next video to generate longer videos.
[…]
40 frames 0.142 0.215 0.148 193.98
80 frames 0.145 0.203 0.139 210.86
120 frames 0.131 0.197 0.121 256.32
160 frames 0.123 0.179 0.101 323.72
@@ 4.5 Ablation Study
Influence of Beat Information. With the emergence of large foundation models capable of generating music representation, many current approaches rely exclusively on these models to encode music, often overlooking information from traditional audio analysis tools. In our experiment, we explore the impact of incorporating beat information into our dance video generation process. The findings, presented in Tab. 6, show that while the inclusion of beat information does not substantially alter the overall quality of the generated videos, it does enhance the performance in terms of video-audio synchronization and motion-beat alignment. Specifically, the model that integrates beat information outperforms the one without in all assessed metrics, underscoring the beneficial effects of beat information.
Arbitrary-Length Video Generation. Given the specialty of our task, our model theoretically can generate videos corresponding to any length of music. By using the final frame of one video sequence as the initial frame of the subsequent sequence, we can create videos of arbitrary length. We conduct experiments with four varying lengths of videos generated sequentially using our model; the results are presented in Tab. 7. As video length increases, quality tends to drop rapidly. This decline is primarily because the final frames generated by the model lack the fidelity of actual video frames, becoming more apparent with longer videos. Additionally, alignment scores diminish due to this quality degradation. For instance, employing a pose estimation algorithm to determine keypoints for calculating the 2D-MM Alignment score becomes less accurate with lower video quality, adversely affecting the alignment score.

### v3

@@ Abstract
Generating dance from music is crucial for advancing automated choreography. Current methods typically produce skeleton keypoint sequences instead of dance videos and lack the capability to make specific individuals dance, which reduces their real-world applicability. These methods also require precise keypoint annotations, complicating data collection and limiting the use of self-collected video datasets. To overcome these challenges, we introduce a novel task: generating dance videos directly from images of individuals guided by music. This task enables the dance generation of specific individuals without requiring keypoint annotations, making it more versatile and applicable to various situations. Our solution, the Dance Any Beat Diffusion model (DabFusion), utilizes a reference image and a music piece to generate dance videos featuring various dance types and choreographies. The music is analyzed by our specially designed music encoder, which identifies essential features including dance style, movement, and rhythm. DabFusion excels in generating dance videos not only for individuals in the training dataset but also for any previously unseen person. This versatility stems from its approach of generating latent optical flow, which contains all necessary motion information to animate any person in the image. We evaluate DabFusion’s performance using the AIST++ dataset, focusing on video quality, audio-video synchronization, and motion-music alignment. We propose a 2D Motion-Music Alignment Score (2D-MM Align), which builds on the Beat Alignment Score to more effectively evaluate motion-music alignment for this new task. Experiments show that our DabFusion establishes a solid baseline for this innovative task. Video results can be found on our project page: https://DabFusion.github.io.
Figure 1: We introduce DabFusion, a diffusion-based framework designed to generate videos of individuals dancing, utilizing a music input and an initial reference image as the conditions for video generation.
@@ 1 Introduction
Music-to-dance generation not only revolutionizes choreography by automating complex routines but also enhances performance training, enriches interactive gaming experiences, and opens new avenues for virtual reality entertainment, offering profound benefits across multiple real-world applications. Existing methods [21, 22, 39, 27] focus on generating keypoint sequences, which, despite being effective, often produce outputs that are less intuitive and do not typically support the animation of specific individuals, which limits their practical utility in personalized applications. Ren et al. [29] attempt to advance the field, but their approach still begins by generating keypoint sequences, which are then used to synthesize videos of specific individuals. This method requires an additional motion transfer model and a video of the target person, complicating the pipeline and limiting practical applications, as obtaining a video of the target person can be challenging in real-world scenarios. Another challenge with existing techniques is their dependency on precisely annotated keypoint data. This requirement not only complicates the data collection process but also restricts the usability of these methods with self-collected or less structured video datasets. To address these limitations, we introduce a new and innovative task: generating dance videos directly from images of individuals, guided by accompanying music. By utilizing images of individuals, this new task facilitates the generation of personalized dance sequences that can feature specific people. Furthermore, it eliminates the need of keypoint annotations, thereby increasing the versatility and applicability of the task across a variety of scenarios. Our solution, named the Dance Any Beat Diffusion model (DabFusion), takes a reference image paired with a selected piece of music as input to create dance videos featuring various dance types and choreographies. For extracting information from music, we develop a novel music encoder that captures essential features including dance style, movement, and rhythm, which are critical for this task. To accurately extract dance styles, we fine-tune the CLAP [51] model, a text-audio pretraining foundation model, on the AIST++ [22] dataset, which serves as the training dataset for our task. For movement information, we fine-tune the Wav2CLIP [50] model, a visual-audio pretraining foundation model, on the AIST++ dataset. The synchronization of dance movements with the music’s rhythm is essential for realistic dance generation. To achieve this, we employ Librosa [23], a tool for audio signal analysis, to extract beat information from the music. By integrating these elements, our music encoder can generate a comprehensive representation of the music.
@@
DabFusion effectively generates dance videos not just for individuals included in the AIST++ dataset, but also for people it has never encountered before. As shown in Fig. 1, we animate previously unseen individuals by merging their images with a background from the AIST++ dataset, bringing them to life through dance. DabFusion can choreograph dance sequences for previously unseen individuals to various music pieces, producing distinctly different videos. This diversity in output highlights the robust information extraction capability of our music encoder. When generating dance videos using the AIST++ dataset, as illustrated in Fig. 2, DabFusion excels in producing varied dance styles with different dancers from multiple perspectives and diverse initial poses. The capability of DabFusion to generate various dance forms enables it to serve a broad audience, ranging from professional choreographers seeking inspiration to casual users wanting to see themselves or others dance to their favorite tunes. The versatility of DabFusion arises from its method of generating latent optical flow, which captures all essential motion information needed to animate any person in the image. This technique is widely adopted in the field of image-to-video (I2V) generation [46, 31, 8, 4, 47, 30, 37, 56]. Inspired by recent I2V works [25, 36, 49], our model’s training process is structured into two phases. First, a latent flow auto-encoder is trained unsupervisedly to estimate the latent optical flow between reference and driving frames in a video, aiding in warping the reference frame to generate movement. Then, a U-Net based diffusion model generates latent flows guided by music and a starting image. We conduct a thorough evaluation of DabFusion’s performance using the AIST++ dataset, with particular focus on video quality, audio-video synchronization, and the alignment between motion and music. To more effectively assess motion-music alignment—a crucial aspect of this new task—we introduce a 2D motion-music alignment (2D-MM Align) score, inspired by the Beat Alignment Score [22] for 3D motion-music correlation. This metric evaluates synchronization between motion and music in 2D scenarios, demonstrating the effectiveness of our approach in producing rhythmically aligned dance videos. Our experimental results demonstrate that DabFusion establishes a robust baseline for this innovative task, indicating its potential as a pioneering tool in the field of dance video synthesis.
@@ 2 Related Works
Generating dance sequences from music uniquely intersects motion synthesis [1, 2, 9, 11] and music interpretation [50, 23, 51], aiming to create choreographed movements synchronized with input music. This extends beyond traditional motion synthesis, as choreographed movements are complex to animate. Early research [35, 57, 6] focused on producing 2D dance sequences due to the availability of online dance videos and advances in 2D human pose estimation [3]. However, 2D predictions lack expressiveness and applicability, prompting a shift to 3D dance generation. Recent methods leverage LSTMs [41, 16, 52], GANs [20, 40, 7, 17], and Transformers [13, 22, 39, 27, 54] for 3D motion generation. The AIST++ dataset [22], a rich compilation of 3D motion data, has significantly advanced the field. Notable models include FACT [22], a full-attention-based cross-modal transformer using sequence-to-sequence learning for lifelike 3D dance sequences, and Bailando [39], which combines a pose VQ-VAE with a Motion GPT for temporal coherence via actor-critic learning. While existing research focuses on generating motion sequences from music, Ren et al. [29] synthesized videos using these sequences. Our work bypasses intermediary joint sequence generation, directly creating dance videos from music.
Figure 3: Overview of DabFusion. Given a reference image $x_{0}$ with dimensions $H_{x}\times W_{x}\times 3$ and a piece of music $m$. DabFusion incorporates noise input along with image embedding $z_{0}$ which has dimensions $H_{z}\times W_{z}\times C_{z}$ and music embedding $e$ as conditions. Following the denoising stage of the diffusion model, we obtain $h_{1}^{N}$ which has dimensions $H_{z}\times W_{z}\times 3\times N$ , comprising a concatenated sequence of latent flow and corresponding occlusion maps. $h_{1}^{N}$ is utilized to transform $z_{0}$ into a new sequence of latent maps, denoted as $\tilde{z}_{1}^{N}$, which is subsequently decoded to produce an image sequence.
@@ 3.2 Music Encoding
Extracting meaningful information from music is essential for generating dance videos, where dance type, dancer movement, and rhythm play critical roles. For dance type information, we utilize the CLAP model, a pipeline for contrastive language-audio pretraining that develops audio representations by integrating audio data with natural language descriptions. We fine-tune CLAP to enhance its capture of dance style information. Specifically, we freeze the weights of the audio and text encoders in CLAP and introduce multi-layer perceptrons as adapter layers. These layers are trained using data from the AIST++ dataset. For movement information, we employ Wav2CLIP, initially trained on audio-visual datasets. We apply a similar fine-tuning approach as with CLAP, using the AIST++ dataset to adapt Wav2CLIP for our specific requirements. Regarding rhythm, which is crucial for aligning dance movements with the music beat, we use Librosa, a tool for audio signal analysis, to extract beat information. Given a piece of music $m$, we pass it through fine-tuned CLAP, fine-tuned Wav2CLIP, and Librosa to extract different features. Specifically, we obtain the music representation $e_{c}$ from CLAP, $e_{w}$ from Wav2CLIP, and beat information $e_{b}$ from Librosa. We then concatenate $e_{c}$, $e_{w}$, and $e_{b}$ to form a comprehensive musical representation $e$, which is specifically tailored for generating dance videos.
@@ 4.1 Dataset and Implementation Details
DabFusion was trained using two A6000 GPUs. We organize our dataset by music pieces, allocating 10,564 videos and 50 music tracks to the training set, and 2,106 videos with 10 tracks to the test set. Each video is segmented into frames, resized to $128\times 128$ pixels. When fine-tuning CLAP and Wav2CLIP, the adapter layers are composed of two MLP layers, each with a hidden size of 512. These adapter layers are trained over 100 epochs. In the latent flow estimation phase, we adopt the architecture from [15] for both our image encoder and decoder. The flow predictor follows [36]. The model is trained over 150 epochs with a batch size of 100, using the Adam optimizer [18], starting with a learning rate of $2\times 10^{-4}$, decreased by a factor of 0.1 after epochs 60, 90, and 120. For the denoising model’s training, we employ a conditional 3D U-Net architecture from [10], featuring four down-sampling and up-sampling 3D convolutional blocks. The time step $t$ is encoded by the sinusoidal position embedding [43]. The training spanned 250 epochs, starting with the same initial learning rate as the first stage, and adjusted at epochs 100, 150, and 200. We set $T$ = 1000 and apply a cosine noise schedule [26] and dynamic thresholding [33] at the 90% during sampling.
Table 1: Quantitative results between DabFusion and MM-Diffusion.
[…]
MM-Diffusion 180.31 0.025 34.81 0.951
DabFusion (ours) 193.98 0.033 26.88 0.959
@@ 4.3 Result Analysis
Video Quality Evaluation. To evaluate the quality of the videos generated by our DabFusion, we use the music and the first frame from the test set, resulting in 2,106 generated videos. We visually present outputs from our model alongside ground truth in Fig. 5. As shown in Fig. 5, the quality of the videos generated by our model closely approximates that of real videos. Given the novelty of our task and the lack of directly comparable previous works, we benchmark our DabFusion against MM-Diffusion [32], the leading model that in unconditional video generation. The AIST++ dataset uniquely utilizes nine cameras ranging from C01 to C09 to record dances from multiple perspectives. DabFusion is trained on the whole dataset and MM-Diffusion is trained on data from only the C01 camera which is in front of the dancer. We randomly generate an equivalent number of videos with MM-Diffusion for comparison. A quantitative comparison between our models and MM-Diffusion is provided in Tab. 1. The results reveal that the quality of videos produced by our models closely matches that of MM-Diffusion. Given that our model can generate videos from various angles, producing videos from specific angles such as C02, which is situated in the upper right of the dance space, proves more challenging than from C01. Despite this, DabFusion demonstrates exceptional performance in the quality of the generated videos.
@@
To further assess the impact of camera angles on video quality, we analyze four specific angles. These are summarized in Tab. 2, illustrating the quality of videos generated from different perspectives. Our models show better performance in terms of FVD and SSIM scores, particularly with the C01 camera position, indicating higher video quality. Our analysis reveals that videos generated from cameras directly facing the dancer, such as C01 and C03, are generally of higher quality compared to those from angled cameras like C02. We also examine the effect of camera distance by comparing videos from C01 and C09, which have the same angle but differ in distance from the dancer. Results indicate that cameras positioned further away tend to produce better outcomes.
@@
Figure 7: Failure cases for DabFusion in different scenarios. On the left, a person with unclear joint delineations leads DabFusion to produce low-quality video output. On the right, the initial crouching pose, uncommon in Ballet Jazz dance, results in repetitive frames.
@@ 4.5 Ablation Study
Influence of Beat Information. With the rise of large models that generate music representations, many current methods rely solely on these models, often neglecting traditional audio analysis tools. In our experiment, we test the impact of including beat information in our dance video generation. The findings, presented in Tab. 6, reveal that while beat information does not greatly change the overall video quality, it improves synchronization between video and audio, and alignment between motion and beat. Specifically, our model that incorporates beat information performs better across all metrics, highlighting the value of including beat information.
5 Discussion
Our DabFusion model effectively handles most of the dance generation scenarios; however, it does encounter some limitations in specific situations, as illustrated in Fig. 7. The first limitation occurs when a dancer’s skeleton cannot be clearly distinguished, such as in the left case of Fig. 7. Additionally, challenges arise when the dancer’s starting pose is atypical for the dance style being simulated, as shown in the right case of Fig. 7. This will lead DabFusion to generate repetitive frames. Moreover, DabFusion is currently limited to creating videos of a fixed length. While chaining sequences together using the final frame of one video as the starting frame for the next can theoretically extend video length, our tests show that this degrades video quality. This deterioration is due to the final frames lacking the fidelity of typical video frames. Future research could improve the quality of longer videos and enable choreography of new dancers without pre-segmentation, making the process simpler and more versatile.